At a certain temperature, the equilibrium constant, ?c, for this reaction is 53.3.

H2(g)+I2(g)↽−−⇀2HI(g)?c=53.3

At this temperature, 0.500 mol H2 and 0.500 mol I2 were placed in a 1.00 L container to react. What concentration of HI is present at equilibrium?
Answer #1

[12]= opeo mal = 0.5m [+] = 0,5 mol = 0.5m 190) = 2456) 0.5 0.5 -a -a +23 Initial +za 2 Change Eq 0.5-2 0.5-2 za Kc= [HI] 531-653.3) + 6533)Y—4(49.3) (13.325) 23 & (493) 533 2840.89–2629.69 98.6 az 533 1. 14:53 986 a = 53.3-14:53 98.5 = 0.39 a=0.39 N